# Axions on a Hyperbolic Ride: Geometric Suppression of CMB Isocurvature and a Blue-Tilted Spectrum

### v2

@@ Abstract
CMB limits on cold-dark-matter isocurvature are often interpreted as excluding the simultaneous realization of high-scale inflation and large QCD axion decay constants in pre-inflationary Peccei–Quinn (PQ) scenarios. We show that this conclusion can be evaded by exploiting field-space geometry. For a minimal complex PQ scalar with a $U(1)$-symmetric potential and a nonlinear sigma-model kinetic term $d\sigma^{2}=dR^{2}+f^{2}(R)\,d\theta^{2}$, the observable axion fluctuation is $\delta\theta\sim H_{\rm inf}/f(R)$, so an enhanced effective decay constant $f(R)$ suppresses isocurvature without explicit PQ breaking, extreme radial displacements, or additional couplings. Specializing to a hyperbolic metric $f(R)\propto\sinh(R/L)$ with curvature scale $L$, we find that for $R\gtrsim L$ the exponential geometric-factor $f(R)$ strongly suppresses CMB-scale isocurvature while producing a characteristic blue-tilted spectrum. The same geometry induces a time-dependent $\mathcal{O}(H_{\rm inf})$ effective mass for the canonical axial mode even during radial slow-roll, which controls the scale-dependence of isocurvature. As a result, inflationary Hubble scales as large as $H_{\rm inf}\sim 10^{13}\,\mathrm{GeV}$ can be compatible with $f_{a}\sim 10^{14}$$10^{16}\,\mathrm{GeV}$, reopening parameter space usually regarded as excluded. We present numerical benchmarks and a semi-analytic template that relates the scale-dependence of isocurvature to the geometric lever arm $R/L$, providing a direct phenomenological probe on PQ field-space geometry.
@@ I Introduction
In this work we present a minimal, $U(1)$ conserving alternative based on field-space geometry. Curved internal field-space metrics are well known to generate effective masses and characteristic scale-dependence for entropic/isocurvature modes in covariant sigma-model systems [8, 46, 14, 35, 24]. In this Letter, we apply this mechanism to a minimal PQ field (in particular, a quartic PQ potential with spontaneously-broken $U(1)_{\rm PQ}$), and show that a hyperbolic field-space metric exponentially suppresses the observable angular fluctuation without explicit PQ breaking. This dynamically suppresses CMB-scale axion isocurvature, while the evolving mass produces a distinctive blue-tilted isocurvature spectrum on smaller scales. As a result, viable parameter space with $H_{\rm inf}\sim 10^{13}\,{\rm GeV}$ and large $f_{a}$ (e.g. $\lesssim 10^{16}\,{\rm GeV}$) can be reopened without requiring the extreme displacements $R_{\rm inf}/f_{a}\sim 10^{4}$$10^{5}$ or tiny quartic coefficients $\lesssim\mathcal{O}(10^{-16})$. We further provide a compact phenomenological template relating the isocurvature spectral index to the field-space curvature scale, thereby promoting measurements of isocurvature scale-dependence to a direct probe of PQ field-space geometry and enabling targeted searches in future analyses.
@@ II The standard QCD axion isocurvature problem
Figure 1: Illustrative constraints in the $(H_{\rm inf},f_{a})$ plane for pre-inflationary QCD axion DM from misalignment. For the standard flat-metric case $f_{\rm eff}=f_{a}$, the isocurvature bound restricts QCD axion to the blue shaded region. Gray bands indicate additional excluded regions. In this work, we show that the white region can be reopened using geometric suppression. (Redrawn with inspiration from [29].)
[…]
$\frac{f_{a}}{2.04\times 10^{18}{\rm GeV}}>\left(\frac{H_{\rm inf}}{10^{10}{\rm GeV}}\frac{1}{\kappa}\right)^{12/5}.$ (6)
$T_{\rm inst.RH}$ is the instantaneous reheating temperature and the orange solid line marks the upper boundary of the region for PQ restoration post-inflation. In flat-metric PQ models, realizing simultaneously $H_{\rm inf}\simeq 10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{16}\,{\rm GeV}$ typically forces $\kappa\gtrsim 10^{4}$, often into a regime where the spectator/EFT description is difficult to control. This motivates alternative suppression mechanisms.
@@ III Hyperbolic PQ field-space and geometric suppression
See Supplemental Material at [52] for derivations in this section. Details and derivations in this section are deferred to the Supplemental Material. We work in the vanishing-charge background $\dot{\theta}_{0}=0$, for which the homogeneous evolution is purely radial,
@@
with curvature scale $L>0$. See Appendix A for a UV-motivated example that realizes this metric. Then
@@ IV Benchmark models and isocurvature spectrum
In the harmonic regime, the total matter isocurvature spectrum is
$\Delta_{S}^{2}(k)\simeq\frac{4\,r_{a}^{2}}{\theta_{i}^{\,2}}\,\Delta_{\theta}^{2}(k),$ (21)
evaluated at the end of inflation $N_{\rm inf}$. For QCD axion DM from misalignment, we eliminate $\theta_{i}$ using Eq. (5).
@@ IV.1 Results
Fig. 2 shows the mass history for B1. The radial field slow-rolls toward its vacuum, while the hyperbolic metric induces a geometric axial mass. Initially $m_{\psi}^{2}\sim\mathcal{O}(H_{\rm inf}^{2})$ while $m^{2}_{R}\ll H_{\rm inf}^{2}$, so modes exiting early are damped. As $R_{0}$ decreases and $\dot{R}_{0}$ redshifts, $m_{\psi}^{2}/H_{\rm inf}^{2}$ falls and the damping weakens for later-exiting modes. In the regime $\delta=\xi(m_{R}/H_{\rm inf})^{2}\lesssim 1$, the dominant behavior is $m_{\psi}^{2}\simeq\xi\,m_{R}^{2}$, i.e. the angular mode is enhanced relative to the flat case by the curvature lever arm $\xi=R/L$, while for $\xi m_{R}^{2}\gtrsim 1$ the negative curvature term $\propto-\dot{R}^{2}$ reduces the effective axial mass.
@@
Fig. 3 shows the resulting conserved superhorizon $\Delta_{\mathcal{S}}^{2}(k)$. We take $k/(a_{i}H_{\rm inf})=10$ to correspond to the CMB pivot scale $k_{\ast}\simeq 0.002,{\rm Mpc}^{-1}$. Power is strongly suppressed on the largest CMB scales because those modes exit while $f(R)$ is exponentially enhanced, so that the observable fluctuation $\delta\theta=\delta\psi/f(R)$ is strongly suppressed at horizon exit. As $R$ slow-rolls during the subsequent radial evolution, $f(R(t_{k}))$ decreases and smaller scales are progressively less suppressed. The outcome is a blue-tilted isocurvature spectrum with running over the range of modes that probe the evolution of $R$ (and hence $f$), approaching a quasi-plateau once the angular mode becomes effectively light. Equivalently, this behavior can be described in terms of the time-dependent effective mass $m_{\psi}^{2}(t)$ during radial slow-roll, which governs the superhorizon evolution of $\delta\psi$ while leaving the ratio $\delta\psi/f(R)=\delta\theta$ nearly conserved in the slow-roll regime outside the horizon.
The benchmarks are chosen such that $\Delta_{\mathcal{S}}^{2}$ at CMB scales lies safely below current bounds, while smaller scales can be enhanced (with isocurvature-to-adiabatic ratio $\sim\mathcal{O}(1)$ near $k\sim 0.1\,{\rm Mpc}^{-1}$, consistent with current limits [18, 3, 9]). This small-scale enhancement is a characteristic feature of the mechanism, and can be modulated by taking $r_{a}<1$, varying $(f_{a},H_{\rm inf})$ within allowed bands, or modifying the radial trajectory (and thus $m_{\psi}^{2}(t)$) through microscopic parameters.
A key result is that hyperbolic geometry can reopen parameter space that is excluded in the flat-metric case. In our benchmarks, we focus on $H_{\rm inf}=10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{14}\text{--}10^{16}\,{\rm GeV}$, as it lies near current upper limits and is the most demanding regime for isocurvature. In contrast, for $f(R)=R$ and comparable spectator trajectories, satisfying the CMB bound at $H_{\rm inf}=10^{13}\,{\rm GeV}$ typically requires extreme radial excursion ($R_{i}/f_{a}\sim 10^{4}\text{--}10^{5}$), often implying super-Planckian displacements.
@@
where $k_{m}$ is a matching scale at the onset of the geometric regime and $\xi_{m}\equiv R_{m}/L$ controls the shape. For the benchmark scenarios, a convenient prescription is to fix $k_{m}$ by $n_{\rm iso}(k_{m})=2$, equivalently $m_{\psi}^{2}(k_{m})/H_{\rm inf}^{2}=5/4$, so that for $k\gtrsim k_{m}$ the entire running is governed by the single curvature parameter $\xi_{m}$, which labels the family of trajectories.
@@ V Conclusion
We have shown that the intrinsic geometry of the PQ field manifold can provide a minimal and symmetry-preserving solution to the standard axion isocurvature problem. For a hyperbolic PQ target-space metric, the angular fluctuation acquires a geometric, time-dependent effective mass during inflation, even in the absence of explicit PQ breaking in the axion potential. The same geometry exponentially enhances $f(R)$, thus suppressing CMB-scale isocurvature while yielding a new mechanism to generate blue-tilted spectrum even during radial slow-roll. In benchmark realizations with $H_{\rm inf}=10^{13}\,{\rm GeV}$, the mechanism reopens parameter space at large $f_{a}$ that is typically excluded in flat field-space.
Beyond enlarging the viable QCD-axion window for high-scale inflation, our results highlight a broader point: measurements of isocurvature scale-dependence can directly probe field-space curvature. We provided a compact phenomenological template that ties the running of the isocurvature tilt to a single curvature-controlled parameter, enabling model-agnostic tests of the mechanism and straightforward incorporation into future analyses. Several broad directions are motivated by this perspective. On the theory side, it will be important to embed controlled UV realizations of hyperbolic metrics to the PQ model, and extend the framework to more general target spaces (non-constant curvature and/or non-diagonal metrics). On the dynamics side, nontrivial background motion will be particularly interesting (such as $\dot{\theta}\neq 0$ or post-inflationary evolution). Hyperbolic geometry can naturally support large angular velocities, making contact with rotating-field mechanisms such as Affleck-Dine dynamics in baryogenesis ([2]), axiogenesis [22] and kinetic-misalignment in [21], and more generally with additional signatures beyond the power spectrum [51, 28]. Finally, the generic enhancement of power on sub-CMB scales motivates systematic confrontation with small-scale probes, including Lyman-$\alpha$ forest [9], UV luminosity functions of high-$z$ galaxies [57], 21-cm measurements [45], and strong-lensing substructure [23, 56], thereby enabling targeted searches for scale-dependent isocurvature as a window onto PQ field-space geometry.

### v3

@@ Abstract
CMB limits on cold-dark-matter isocurvature are often interpreted as excluding the simultaneous realization of high-scale inflation and large QCD axion decay constants in pre-inflationary Peccei–Quinn (PQ) scenarios. This conclusion can be evaded by exploiting field-space geometry. For a minimal complex PQ scalar with a $U(1)$-symmetric potential and a nonlinear sigma-model kinetic term $d\sigma^{2}=dR^{2}+f^{2}(R)\,d\theta^{2}$, the observable axion fluctuation is $\delta\theta\sim H_{\rm inf}/f(R)$, so an enhanced effective decay constant $f(R)$ suppresses isocurvature without explicit PQ breaking, extreme radial displacements, or additional couplings. We specialize to a hyperbolic metric $f(R)\propto\sinh(R/L)$ with curvature scale $L$. The same geometry also induces a time-dependent $\mathcal{O}(H_{\rm inf})$ effective mass for the canonical axial mode during radial slow-roll, and fixing the tilt and running of isocurvature. Thus, CMB-scale isocurvature is suppressed while a characteristic blue-tilted spectrum is generated. As a result, inflationary Hubble scales as large as $H_{\rm inf}\sim 10^{13}\,\mathrm{GeV}$ can be compatible with $f_{a}\sim 10^{14}$$10^{16}\,\mathrm{GeV}$, reopening parameter space usually regarded as excluded. We present ‘observable’ benchmarks and a semi-analytic template that relates the scale-dependence of isocurvature to the geometric lever arm $R/L$, providing a direct phenomenological probe on PQ field-space geometry.
@@ I Introduction
In this work we present a minimal, $U(1)$ conserving alternative based on field-space geometry. Curved internal field-space metrics are well known to generate effective masses and characteristic scale-dependence for entropic/isocurvature modes in covariant sigma-model systems [8, 47, 14, 35, 24, 43]. In this Letter, we apply this mechanism to the QCD axion, modeled by a minimal PQ complex scalar field with a quartic potential and a spontaneously broken $U(1)_{\rm PQ}$. A hyperbolic field-space metric then exponentially suppresses the observable angular fluctuation without explicit PQ breaking, while the evolving axion mass produces a distinctive blue-tilted isocurvature spectrum on smaller scales. As a result, viable parameter space with $H_{\rm inf}\sim 10^{13}\,{\rm GeV}$ and large $f_{a}$ (e.g. $\lesssim 10^{16}\,{\rm GeV}$) can be reopened without requiring the extreme displacements $R_{\rm inf}/f_{a}\sim 10^{4}$$10^{5}$ or tiny quartic coefficients $\lesssim\mathcal{O}(10^{-16})$. The effective radial mass and the geometric lever arm control the ‘detectability’ of these isocurvatures. Thus, we further provide a compact phenomenological template relating the isocurvature spectral index to the field-space curvature scale, thereby promoting measurements of isocurvature scale-dependence to a direct probe of PQ field-space geometry.
@@ II The standard QCD axion isocurvature problem
Figure 1: Illustrative constraints in the $(H_{\rm inf},f_{a})$ plane for pre-inflationary QCD axion DM from misalignment. For the standard flat-metric case $f_{\rm eff}=f_{a}$, the isocurvature bound restricts QCD axion to the blue shaded region. Gray bands indicate additional excluded regions. We focus on how the white region can be reopened using geometric suppression. (Redrawn with inspiration from [29].)
Fig. 1 displays the resulting tension in the $(H_{\rm inf},f_{a})$ plane for the usual QCD axion. If the inflationary effective decay constant is enhanced to $f_{\rm eff}=\kappa f_{a}$, the bound is relaxed accordingly. $T_{\rm inst.RH}$ is the instantaneous reheating temperature and the orange solid line marks the upper boundary of the region for PQ restoration post-inflation. In flat-metric PQ models, realizing simultaneously $H_{\rm inf}\simeq 10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{16}\,{\rm GeV}$ typically forces $\kappa\gtrsim 10^{4}$, often into a regime where the spectator/EFT description is difficult to control.
@@ III Hyperbolic PQ field-space and geometric suppression
See Supplemental Material at [53] for derivations in this section. We work in the vanishing-charge background $\dot{\theta}_{0}=0$, for which the homogeneous evolution is purely radial,
@@
with curvature scale $L>0$. See Appendix A for a UV-motivated example that realizes this metric through a Kähler manifold. Then
@@ IV Benchmark models and isocurvature spectrum
In the harmonic regime, the total matter isocurvature spectrum is given by Eq. (3) evaluated at the end of inflation $N_{\rm inf}$. For QCD axion DM from misalignment, we eliminate $\theta_{i}$ using Eq. (5).
@@ IV.1 Results
Fig. 2 shows the mass history for B1. The radial field slow-rolls toward its vacuum, while the hyperbolic metric induces a geometric axial mass. Initially $m^{2}_{R}\lesssim\mathcal{O}(0.1)H_{\rm inf}^{2}$ while $m_{\psi}^{2}\sim\mathcal{O}(H_{\rm inf}^{2})$, so modes exiting early are damped. As $R_{0}$ decreases and $\dot{R}_{0}$ redshifts, $m_{\psi}^{2}/H_{\rm inf}^{2}$ falls and the damping weakens for later-exiting modes. In the regime $\delta=\xi(m_{R}/H_{\rm inf})^{2}\lesssim 1$, the dominant behavior is $m_{\psi}^{2}\simeq\xi\,m_{R}^{2}$, i.e. the angular mode is enhanced relative to the flat case by the curvature lever arm $\xi=R/L$, while for $\xi m_{R}^{2}\gtrsim 1$ the negative curvature term $\propto-\dot{R}^{2}$ reduces the effective axial mass.
@@
Fig. 3 shows the resulting conserved superhorizon $\Delta_{\mathcal{S}}^{2}(k)$. We take $k/(a_{i}H_{\rm inf})=10$ to correspond to the CMB pivot scale $k_{\ast}\simeq 0.002,{\rm Mpc}^{-1}$. Power is strongly suppressed on the largest CMB scales because those modes exit while $f(R)$ is exponentially enhanced, so that the observable fluctuation $\delta\theta=\delta\psi/f(R)$ is strongly suppressed at horizon exit [43]. As $R$ slow-rolls during the radial evolution, $f(R(t_{k}))$ decreases and smaller scales are progressively less suppressed. The outcome is a blue-tilted isocurvature spectrum with running over the range of modes that probe the evolution of $R$, approaching a quasi-plateau once the angular mode becomes light. Equivalently and rather more accurately, this behavior can be described in terms of the time-dependent effective mass $m_{\psi}^{2}(t)$ during radial slow-roll, which governs evolution of $\delta\psi$ while leaving the ratio $\delta\psi/f(R)=\delta\theta$ nearly conserved in the slow-roll regime outside the horizon.
IV.2 Detectability of isocurvature signal
For $R/L\gg 1$, the hyperbolic metric implies $\Delta_{\theta}^{2}\sim\left(H_{\rm inf}/L\right)^{2}e^{-2R/L}$, so that for $R/L\gg 10$, the isocurvature amplitude is negligible and phenomenologically irrelevant. For representative values $H_{\rm inf}\lesssim L$, near-term observability therefore requires $R\lesssim\mathcal{O}(10)\,L$.
Within this detectable window, the spectral morphology is controlled by radial dynamics. If the radial mode is strongly overdamped, $m_{R}^{2}\lesssim\mathcal{O}(0.01)H_{\rm inf}^{2}$, its evolution over the observable e-folds is negligible. For $R/L\lesssim\mathcal{O}(10)$, $m_{\psi}^{2}\lesssim\mathcal{O}(0.1)H_{\rm inf}^{2}$, yielding an approximately scale-invariant spectrum with negligible running. Such a signal may remain detectable at CMB scales but does not generate distinctive small-scale structure. In contrast, for mild slow-roll with $m_{R}^{2}\lesssim\mathcal{O}(0.1)H_{\rm inf}^{2}$, the radial field evolves by order unity and the geometry-induced mass satisfies $m_{\psi}^{2}\sim H_{\rm inf}^{2}$. The resulting axion isocurvature spectrum is then significantly blue-tilted, suppressed on CMB scales yet enhanced on small scales, with potentially observable small-scale consequences. The same radial evolution also induces measurable running, providing a characteristic signature of the geometric mechanism.
The benchmarks are therefore chosen with $R/L\sim 10$ such that $\Delta_{\mathcal{S}}^{2}$ at CMB scales lies safely below current bounds, while smaller scales can be enhanced (with isocurvature-to-adiabatic ratio $\sim\mathcal{O}(1)$ near $k\sim 0.1\,{\rm Mpc}^{-1}$, consistent with current limits [18, 3, 9]).
The above result highlight that hyperbolic geometry can reopen parameter space that is excluded in the flat-metric case. In our benchmarks, we focussed on $H_{\rm inf}=10^{13}\,{\rm GeV}$ and $f_{a}\sim 10^{14}\text{--}10^{16}\,{\rm GeV}$, as it lies near current upper limits and is the most constrained regime for isocurvature. In contrast, for $f(R)=R$, satisfying the CMB bound at $H_{\rm inf}=10^{13}\,{\rm GeV}$ typically requires extreme radial excursion ($R_{i}/f_{a}\sim 10^{4}\text{--}10^{5}$), implying super-Planckian displacements.
@@ IV.2 Detectability of isocurvature signal
where $k_{m}$ is a matching scale at the onset of the geometric regime and $\xi_{m}\equiv R_{m}/L$ controls the shape. For the benchmark scenarios, a convenient prescription is to fix $k_{m}$ by $n_{\rm iso}(k_{m})=2$, equivalently $m_{\psi}^{2}(k_{m})/H_{\rm inf}^{2}=5/4$, so that for $k\gtrsim k_{m}$ the entire running is governed by the single parameter $\xi_{m}$, labelling the family of trajectories.
@@ V Conclusion
We have demonstarted that the intrinsic geometry of the PQ field manifold provides a minimal and symmetry-preserving solution for the axion isocurvature problem. The curved (hyperbolic) field-space enhances the effective decay constant and can suppress CMB-scale isocurvature. The angular fluctuation acquires a geometric, time-dependent effective mass during inflation, without any explicit PQ breaking in the axion potential.
The phenomenology is governed by the geometric lever arm $R/L$ and radial dynamics. For moderate $R/L\lesssim\mathcal{O}(10)$, the signal can remain detectable and CMB-compatible while radial slow-roll enhances small-scale power and induces measurable running. If the radial mode is strongly overdamped, the spectrum is nearly scale-invariant with negligible running, whereas very large $R/L$ suppresses the signal altogether. In benchmark realizations with $H_{\rm inf}=10^{13}\,{\rm GeV}$, this mechanism reopens parameter space at large $f_{a}$ that is typically excluded in flat field-space.
Beyond enlarging the viable QCD-axion window at high inflationary scales, our results highlight that isocurvature scale-dependence can directly probe field-space curvature. We provided a compact template linking the running of the isocurvature tilt to a single curvature-controlled parameter, enabling model-agnostic tests of the mechanism and straightforward incorporation into future analyses.
Several broad directions are motivated by this perspective. On the theory side, constructing controlled UV realizations of hyperbolic metrics in PQ models, and extending the framework to more general target spaces (e.g. non-constant curvature or non-diagonal metrics) are important. On the dynamics side, nontrivial background motion such as $\dot{\theta}\neq 0$ or post-inflationary evolution may introduce additional observational constraints, or signatures beyond the power spectrum [52, 28]. Within rotating-field mechanisms such as Affleck-Dine dynamics in baryogenesis ([2]), axiogenesis [22] and kinetic-misalignment in [21], hyperbolic geometry can support large angular velocities. Finally, the predicted enhancement of sub-CMB power motivates systematic confrontation with small-scale probes, including Lyman-$\alpha$ [9], UV luminosity functions [58], 21-cm measurements [46], and strong-lensing substructure [23, 57], offering a direct observational window onto PQ field-space geometry.